OpenStudy (anonymous):

Solve for x: 7 over 8 minus 1 over x equals 3 over 4

OpenStudy (zzr0ck3r):

\[\frac{7}{8}-\frac{1}{x}=\frac{3}{4}\] multiply everything by 8x because 8,x,4 all divide 8x \[8x\frac{7}{8}-8x\frac{1}{x}=8x\frac{3}{4}\] simplify \[7x-8=6x\] can you solve?